OpenStudy (anonymous):

12 + 4m = 24 m = 9 m = 8 m = 3 m = -6

OpenStudy (anonymous):

m=3

OpenStudy (nurali):

12 + 4m = 24 4m = 24-12 4m=12 m=12/4 m=3

OpenStudy (anonymous):

thanks

OpenStudy (nurali):

Anytime.